ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 60, 68, 93 Orman, Gabriel V., Transylvania University of Brasov, Romania On limit theorems and some of their applications From some time past we focus our interest to find new possibilities to characterize the process of generation of the words by generative. systems. To this end we have adopted a stochastic point of view involving Markov chains. Because the sequences of symbols by which the words may be generated are finite, it results that finite Markov chains will be connected to the process. The novelty that we have proposed consists in the organization of the process of generation of the words by considering the equivalence classes of the sequences of symbols by which a word may be generated. In this way our results are available up to an equivalence. Markov dependence characterizations are given and applications in terms of limit theorems are presented. Reference: G. V. Orman, Markov chains connected to the process of generation of the words by a generative system, Yugoslav Journal of Operations Research 4 (1994), No. 2, 167-178.